OpenStudy (anonymous):

find the derivative of f(x) = (3x^2 - 3x + 1)/(2x). Directions, put 2x under each term and simplify. Then take the derivative.

OpenStudy (anonymous):

@satellite73 can u help?

OpenStudy (anonymous):

@onegirl hints has been given in the ques so what will u get after simplifying ?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

after i put 2x over each term f(x) = 3x^2/2x – 3x/2x + 1/2x f(x) 3/2 x + 3/2 + 1/2x f’(x) = 3/2 + 0 - 1/2x^2 f’(x) = 3x^2 – 1

OpenStudy (anonymous):

last step is wrong u forgot to write in denominator?

OpenStudy (anonymous):

what step?

OpenStudy (anonymous):

for f'(x) = 3x^2 - 1?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so 3x^2 -1/2x?

OpenStudy (mertsj):

\[f(x)=\frac{3}{2}x-\frac{3}{2}+\frac{1}{2}x ^{-1}\]

OpenStudy (anonymous):

it will be 3x^2-1/2x^2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what do i do after that?

OpenStudy (anonymous):

it is ur answer

OpenStudy (mertsj):

\[f'(x)=\frac{3}{3}-\frac{1}{2}x ^{-2}\]

OpenStudy (anonymous):

so thats my derivative?

OpenStudy (mertsj):

Whoops. Typo. Should be 3/2 not 3/3

OpenStudy (anonymous):

yeah thats your derivative

OpenStudy (anonymous):

okay ii guess not putting 2x^2 is why i got it wrong

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thx

OpenStudy (mertsj):

You were supposed to put 2x UNDER each term

OpenStudy (anonymous):

yea i did

OpenStudy (anonymous):

@Mertsj so i did it wrong?

OpenStudy (anonymous):

@zepdrix can u help please? did i do it right or wrong?

zepdrix (zepdrix):

Do you understand this step? ~Splitting up the fractions. I wrote those ugly blobs to help show how we'll split it up.